# Charged analogues of a singularity-free anisotropic compact star under linear $f(Q)$- action

### v3

@@ Abstract
This study simulates the characteristics of spherically symmetric, anisotropic compact stellar bodies with electrical charge within the context of the $f(Q)$ theory of gravity. Employing the Krori-Barua metric ansatz (K.D. Krori, J. Barua, J. Phys. A: Math. Gen. 8 (1975) 508) along with a linear form of $f(Q)$ model, viz., $f(Q)=\alpha_{0}+\alpha_{1}Q$, we obtain a tractable set of exact relativistic solutions of the field equations. A specific form of charge $(q=q_{0}r^{3})$ is considered here for the present analysis. Using the data sets of $\rho$ and $p_{r}$ along with the method of curve fitting, we have obtained the best fit equation of state in the model, which is incorporated in the numerical solutions of the TOV equations to determine the maximum mass and radius in this scenario. With increasing charge intensity $(q_{0})$ from 0.0002 to 0.0006, the maximum mass ranges from $2.838-2.869~{}M_{\odot}$, and the corresponding radii range from $12.0039-12.0822~{}Km$. Moreover, the predicted radii of some recently observed pulsars and GW 190814 show that our model also complies with the estimated radii based on the observational results. Our model is found to satisfy all the characteristic features, such as behaviour of matter variables, causality condition, energy constraints and stability criteria, which are pertinent in the context of a stable stellar configuration to emerge as a viable and physically acceptable stellar model in the framework of $f(Q)$ gravity.
@@ 1 Introduction
In 1912, Einstein’s exploration of static gravitational fields led him to a daring proposition. A straightforward application of his principle of equivalence, combined with the fundamental outcomes of special relativity, implied that the gravitational field is characterised by the metric tensor. He hypothesised that this holds true even beyond the static limit and this set him on a three-year quest that concluded in November 1915 with the formulation of the field equations of his General Theory of Relativity (GR). In the later years, GR has solved many persistent mysteries pervading in astrophysics and cosmology. Despite being remarkably successful, GR has faced significant challenges when confronted with astrophysical and cosmological observational results. For instance, numerous distinct cosmological studies suggest that our universe is undergoing an accelerated expansion [1, 2, 3], the phenomenon of dark energy and dark matter, which are essential in galactic dynamics, as well as the accelerated expansion of the universe, etc., are not adequately explained within the framework of GR without introducing exotic forms of matter and energy. Additionally, GR predicts singularities, such as those at the centres of black holes and the Big Bang, where the laws of physics break down. These singularities indicate a need for a more comprehensive theory that can be seamlessly integrated with quantum mechanics for a deeper understanding at the minute level. In this context, modified theories of gravity offer an intricate way to address these challenges by extending the geometrical and dynamical foundations of GR, potentially offering alternative explanations for these enigmatic phenomena and providing a more complete understanding of the universe.
Soon after the advent of GR, Weyl [4] imposed the first modifications on it by introducing higher order invariants to the Einstein-Hilbert action to unify electromagnetism and gravitation. Later, GR was modified by two types of geometrical extensions, viz.,
@@
Such torsion and non-metricity based theories of gravity, which are equivalent to GR, are termed the Teleparallel Equivalent of General Relativity (TEGR) [6, 7] and the Symmetric Teleparallel Equivalent of General Relativity (STEGR) [9, 10]. In the context of flat space-time with torsion, the TEGR is formulated using the tetrads and spin connections as the primary variables. The constraints imposed by the nullification of the curvature and non-metricity tensor restrict the spin connection. This allows for the adoption of the Weitzenböck connection, where all elements of the spin connections are eliminated, leaving only the tetrad components as the fundamental entities. This selection is interpreted as a gauge choice within the framework of TEGR. It is noteworthy that this choice does not influence the physical implications of the theory, as any other selection compatible with the teleparallelism constraint would yield the same action, barring a surface term. On the other hand, in the framework of STEGR, it is the non-metricity that characterises gravity, as opposed to curvature and torsion. Within the constraints of teleparallelism, a particular gauge, known as the coincidence gauge, can be selected in this theory. This gauge selection sets the metric tensor as the sole fundamental variable. An additional extension of the STEGR is the $f(Q)$ gravity theory, which bears a significant resemblance to the $f(R)$ gravity theory [9, 11]. Within the general formalism of $f(Q)$ gravity, Hohmann et al. [12] studied the potential polarisation and the velocity of propagation of gravitational waves in Minkowski space. Subsequently, Soudi et al. [13] showed that gravitational wave polarisation has an important impact on the strong-field nature of theories of gravity. $f(Q)$ theory has been utilised in a wide range of topics, including late-time acceleration [14], black holes [15, 16, 17], bouncing [18], evolution of the growth index in the context of matter fluctuations [19], and $f(Q)$ parametrisation to integrate the observational constraints [20]. A detailed review of $f(Q)$ formalism has been described in Ref. [21]. In addition, several aspects of $f(Q)$ gravity have been described in Refs. [22, 23, 24, 25, 26].
@@
In the astrophysical context, the study of compact objects within the framework of modified gravity is essential for advancing our understanding of fundamental physics. Compact objects such as black holes and neutron stars provide extreme environments where the effects of gravity are exceptionally strong, offering unique testing grounds for theories that extend beyond GR. These objects can reveal discrepancies between observed phenomena and theoretical predictions, shedding light on potential modifications to our current gravitational models. Recently, a substantial volume of data from pulsars and gravitational wave (GW) events has been collected and scrutinised by numerous researchers. This comprehensive analysis has led to precise predictions of several physical parameters pertaining to such celestial bodies. By utilising the Shapiro delay [27] an accurate determination of the mass of millisecond pulsar PSR J0740+6620 resulted to be $2.14^{+0.10}_{-0.09}$. In addition to this, it is postulated that there could be a few compact celestial bodies that possess masses exceeding that of PSR J0740+6620 [27], particularly within the cluster of interacting systems. The occurrence of gravitational wave event, referred to as GW 190814, has indicated that the associated star boasts a mass ranging from 2.5 to 2.67 $M_{\odot}$, with a confidence level of 90%. It is currently unclear whether this component represents an extremely massive compact star or a lightweight black hole. If the former is true, it becomes imperative to introduce a novel concept into the theoretical framework to extend the maximum mass range to account for such elevated mass values. This kind of formalism would be crucial for investigating the internal structure of dense matter beyond saturation density.
@@
The pressure anisotropy within a compact celestial body can be attributed to various factors. One potential cause is the existence of a type 3A superfluid within the solid core of these objects [28]. Due to the nature of fermions, nucleons cannot occupy the same energy state simultaneously, a principle known as the Pauli exclusion principle. Additionally, strong repulsive interactions between nucleons occur within very short ranges. These repulsive forces aid in counteracting gravitational collapse within a neutron star. However, this scenario changes when temperatures are extremely low, leading to the formation of Cooper pairs [29] of nucleons, which essentially behave as bosons. Consequently, at very low temperatures, nucleons exhibit collective behaviour on a large scale. These nucleon condensates, analogous to Helium-3, can flow without any viscosity. Within a neutron star, the high-pressure region effectively raises the critical temperature for Cooper pair formation, allowing for the occurrence of nuclear superfluidity even at temperatures on the order of a billion degrees. There could be three distinct types of superfluids present within the core of a neutron star [30]. Other potential sources of anisotropy in compact objects could include phase transitions [31], pion condensation [32], slow rotation and viscosity [33].
@@
Rosseland [34] was the first to propose that a star could be composed of a substantial number of electrons and positive ions. Given their extremely high kinetic energy, electrons have a higher likelihood of escaping from the star compared to positive ions. As a result, a star can possess a considerable amount of net positive charge. This process persists until the internal electric field prevents further electron escape [35]. The Einstein Field Equations (EFE), when considering an electric field, are crucial for examining the properties of neutron stars, quark stars, gravastars, black holes, and other entities in the realm of relativistic astrophysics. This is because the electric field significantly influences the overall properties of such compact objects. In his study, Ivanov [36] showed that one of the most important prerequisites for obtaining a singularity-free perfect fluid distribution is the presence of non-vanishing net charge. While arguing on the results of Hoyle and Narlikar [37] about the collapse of a large mass of gas into singularity, Bonnor [38] showed that electric repulsion can counterbalance the gravitational collapse of a spherical body to achieve an equilibrium configuration. Stettner [39] showed that a spherical structure with a net surface charge is more stable than a structure with a uniform density. Krasinski [40] demonstrated that preventing the gravitational collapse of spherically distributed matter into a point singularity necessitates the presence of a net charge. Through a generalisation of the Oppenheimer–Volkoff equation, Bekenstein [41] conducted a qualitative assessment of the stability of spheres composed of charged fluid. Over the years, several researchers have addressed the impact of charge on compact stellar structures [42, 43, 44, 45, 46, 47, 48, 49, 50].
@@
Recently, the $f(Q)$ theory has garnered a significant amount of attention from researchers worldwide. In the astrophysical scenario, it has provided new insights into compact stellar structures. In the context of charged configurations, Kaur et al. presented a charged $f(Q)$ solution to study an anisotropic fluid distribution using the Vaidya-Tikekar metric ansatz. Maurya et al. [52] used an interesting approach of gravitational decoupling into the $f(Q)$ theory to study the constraining factors of mass and radii of the lighter component of GW190814 and other self-bound strange stars. Errehymy et al. [53] studied a singularity-free charged strange star model in the framework of $f(Q)$ gravity. Furthermore, Lohakare et al. [54] investigated the effects of gravitational decoupling on the maximum mass and stability of charged strange stars. In addition, several authors have included $f(Q)$ theory in their analysis of a stable stellar structure [55, 56, 57, 58].
@@
In the present manuscript, we have investigated the impact of charge on the structure of a singularity-free compact star model within the framework of linear $f(Q)$ action. The advantage of $f(Q)$ gravity lies in its ability to derive the equations of GR without the necessity of employing the affine connection, thereby enhancing the inertial gravitational acceleration. This intriguing aspect coupled with the charged configuration reveals interesting stellar properties. The rest of the paper is organised in the following manner: Section 2 demonstrates the fundamental ground work of $f(Q)$ theory in the presence of charge. Section 3 addresses the formulation and solution of the Einstein-Maxwell equations in the framework of $f(Q)$ gravity. The analytical expressions of the thermodynamic quantities such as energy density $(\rho)$, radial $(p_{r}$) and tangential $(p_{t})$ pressures are obtained in this section. The effects of $f(Q)$ parameters on the boundary conditions are described in Section 4. Section 5 deals with the physical characteristic features of the model along with the determination of the maximum mass and radius, causality condition and energy conditions. The necessary stability criteria are addressed in Section 6. Finally, we conclude by summarising the major findings of the paper in Section 7.
@@ 2 Fundamentals of $f(Q)$ gravity in the presence of charge
where, $g$ defines the determinant of the fundamental metric tensor $g_{ij}$, i.e., $g=|g_{ij}|$, $f(Q)$ is a function of non-metricity $Q$, $\lambda^{kij}_{l}$ and $\tau^{ij}_{k}$ are two Lagrangian multipliers, $R^{l}_{kij}$ and $T^{k}_{ij}$ are the Riemann tensor and torsion tensor, respectively, and $L_{m}$ and $L_{q}$ describe the Lagrangian density of matter and electromagnetic charge, respectively. Using affine connections $(\Gamma^{k}_{ij})$, the non-metricity is expressed as:
@@
    The final term on the right hand side of Eq. (3) represents the deformation and is expressed as:
@@
Due to the absence of torsion and curvature, the affine connection can be effectively characterised by a tractable set of functions described as:
@@
Now, we obtain the tractable set of gravitational field equations by varying the Einstein-Hilbert action as written in Eq. (1) with respect to the metric tensor $g_{ij}$ in the form:
@@
Moreover, the variation of action with respect to the affine connection yields,

### v4

# Charged analogues of singularity-free anisotropic compact stars under linear $f(Q)$-action

@@ Abstract
This study simulates the characteristics of spherically symmetric, anisotropic compact stellar bodies with electrical charge within the framework of the $f(Q)$ theory of gravity. Employing the Krori-Barua metric ansatz (K.D. Krori, J. Barua, J. Phys. A: Math. Gen. 8 (1975) 508) along with a linear form of $f(Q)$ model, viz., $f(Q)=\alpha_{0}+\alpha_{1}Q$, we obtain a tractable set of exact relativistic solutions of the field equations. A specific form of charge $(q=q_{0}r^{3})$ is considered here for the present analysis. It is noted that the model is valid up to the value of charge intensity $q_{0}\leq 0.0009~{}Km^{-2}$. Beyond this value, the model does not permit physically viable results. We have obtained the best fit equation of state in the model, which is incorporated to solve the TOV equations numerically to determine the mass-radius relation within the parameter space used here. With increasing charge intensity $(q_{0})$ from 0.0002 to 0.0009, the maximum mass ranges from $2.84-2.92~{}M_{\odot}$, and the corresponding radii range from $12.00-12.20~{}Km$. Moreover, the predicted radii of some recently observed pulsars and GW 190814 show that our model also complies with the estimated radii based on the observational results. Our model is found to satisfy all the characteristic features, such as behaviour of matter variables, causality condition, energy constraints and stability criteria, which are pertinent in the context of a stable stellar configuration to emerge as a viable and physically acceptable stellar model in the framework of $f(Q)$ gravity.
@@ 1 Introduction
Einstein’s exploration of static gravitational fields led him to a daring proposition. A straightforward application of his principle of equivalence, combined with the fundamental outcomes of special relativity, implied that the gravitational field is characterised by the metric tensor. He hypothesised that this holds true even beyond the static limit and this set him on a three-year quest that concluded in November 1915 with the formulation of the field equations of his General Theory of Relativity (GR). In the later years, GR has solved many persistent mysteries pervading in astrophysics and cosmology. Despite being remarkably successful, GR has faced significant challenges when confronted with astrophysical and cosmological observational results. For instance, numerous distinct cosmological studies suggest that our universe is undergoing an accelerated expansion [1, 2, 3], the phenomenon of dark energy and dark matter, which are essential in galactic dynamics, as well as the expansion of the size of the universe in accelerated mode, etc., are not adequately explained within the framework of GR without introducing exotic forms of matter and energy. Additionally, GR predicts singularities, such as those at the centres of black holes and the Big Bang, where the laws of physics break down. These singularities indicate a need for a more comprehensive theory that can be seamlessly integrated with quantum mechanics for a deeper understanding at the minute level. In this context, modified theories of gravity offer an intricate way to address these challenges by extending the geometrical and dynamical foundations of GR, potentially offering alternative explanations for these enigmatic phenomena and providing a more complete understanding of the universe.
Soon after the advent of GR, Weyl [4] imposed the first modifications on it by introducing invariants of higher order to the action of Einstein-Hilbert to unify the electromagnetism and gravitation. Later, GR was modified by two types of geometrical extensions, viz.,
@@
Such torsion and non-metricity based theories of gravity, which are equivalent to GR, and, are termed the Teleparallel Equivalent of General Relativity (TEGR) [6, 7] and the Symmetric Teleparallel Equivalent of General Relativity (STEGR) [9, 10]. In the context of flat space-time with torsion, the TEGR is formulated using the tetrads and spin connections as the primary variables. The constraints imposed by the nullification of the curvature and non-metricity tensor restrict the spin connection. This allows for the adoption of the Weitzenböck connection, where all elements of the spin connections are eliminated, leaving only the tetrad components as the fundamental entities. This selection is interpreted as a gauge choice within the framework of TEGR. It is noteworthy that this choice does not influence the physical implications of the theory, as any other selection compatible with the teleparallelism constraint would yield the same action, barring a surface term. On the other hand, in the framework of STEGR, it is the non-metricity that characterises gravity, as opposed to curvature and torsion. Within the constraints of teleparallelism, a particular gauge, known as the coincidence gauge, can be selected in this theory. This gauge selection sets the metric tensor as the sole fundamental variable. An additional extension of the STEGR is the $f(Q)$ gravity theory, which bears a significant resemblance to the $f(R)$ gravity theory [9, 11]. Within the general formalism of $f(Q)$ gravity, Hohmann et al. [12] studied the potential polarisation and the velocity of propagation of gravitational waves in Minkowski space. Subsequently, Soudi et al. [13] showed that gravitational wave polarisation has an important impact on the strong-field nature of theories of gravity. $f(Q)$ theory has been utilised in a wide range of topics, including late-time acceleration [14], black holes [15, 16, 17], bouncing [18], evolution of the growth index in the context of matter fluctuations [19], and $f(Q)$ parametrisation to integrate the observational constraints [20]. A detailed review of $f(Q)$ formalism has been described in Ref. [21]. In addition, several aspects of $f(Q)$ gravity have been described in Refs. [22, 23, 24, 25, 26].
@@
In the astrophysical context, the study of compact objects within the framework of modified gravity is essential for advancing our understanding of fundamental physics. Compact objects such as black holes and neutron stars provide extreme environments where the effects of gravity are exceptionally strong, offering unique testing grounds for theories that extend beyond GR. These objects can reveal discrepancies between observed phenomena and theoretical predictions, shedding light on potential modifications to our current gravitational models. Recently, a large amount of data has been collected from pulsars and gravitational wave (GW) events and scrutinised through numerous research. This comprehensive analysis has led to precise predictions of several physical parameters pertaining to such celestial bodies. By utilising the Shapiro delay [27] an accurate determination of the mass of millisecond pulsar PSR J0740+6620 resulted to be $2.14^{+0.10}_{-0.09}$. In addition to this, it is postulated that there could be a few compact celestial bodies that possess masses exceeding that of PSR J0740+6620 [27], particularly within the cluster of coupled systems. The occurrence of the event of gravitational wave, referred to as GW 190814, has indicated that the associated star boasts a mass ranging from 2.5 to 2.67 $M_{\odot}$, with a confidence level of 90%. It is currently unclear whether it represents a massive compact star or a lightweight black hole. Accordingly, if the former is true, it becomes imperative to introduce a novel concept into the theoretical framework to extend the maximum mass range to account for such elevated mass values. This kind of formalism would be crucial for investigating the internal structure of dense matter beyond saturation density.
@@
The pressure anisotropy within a compact celestial body can be attributed to various factors. One potential cause is the existence of a superfluid of type 3A within the solid core of these objects [28]. Due to the property of fermions, nucleons cannot occupy the same energy state simultaneously. Additionally, strong repulsive interactions between nucleons occur within very short ranges. These repulsive forces aid in counteracting gravitational collapse within a neutron star. However, this scenario changes when temperatures are extremely low, leading to the formation of Cooper pairs [29] of nucleons, which essentially behave as bosons. Consequently, at very low temperatures, nucleons exhibit collective behaviour on a large scale. These assembly of nucleons may condensate to form a state analogous to Helium-3, and, can flow without any viscosity. The high-pressure region, within a neutron star, effectively raises the critical temperature for the formation of Cooper pair, allowing for the occurrence of nuclear superfluidity even at temperatures of the order of a billion degrees. There could be three distinct types of superfluids present within the core of a neutron star [30]. Other potential sources of anisotropy inside compact objects could include the phase transitions [31], pion condensation [32], slow rotation and viscosity [33].
@@
Considering electric charge in the study of compact stars is crucial because intense gravitational fields in such stars can separate charges, leading to a significant electric field. This electric charge can alter the star’s structure, stability, and mass-radius relationship. Including charge in models may reveal new insights about high-energy phenomena, quark matter interactions, and the internal dynamics of stars, especially in extreme conditions found in NS. Rosseland [34] was the first to propose that a star could be composed of a substantial number of electrons and positive ions. Given their extremely high kinetic energy, electrons have a higher likelihood of escaping from the star compared to positive ions. As a result, a star can possess a considerable amount of net positive charge. Such process persists until the internal electric field prevents further electron escape [35]. The Einstein Field Equations (EFE), when considering an electric field, are crucial for examining the viable properties of compact stars such as neutron stars, quark stars, strange stars, gravastars, black holes, and other entities in the realm of relativistic astrophysics. This is because the electric field significantly influences the overall properties of such compact objects. In his study, Ivanov [36] showed that one of the most important prerequisites for obtaining a singularity-free perfect fluid distribution is the presence of non-vanishing net charge. While arguing on the results of Hoyle and Narlikar [37] about the collapse of a large mass of gas into singularity, Bonnor [38] showed that electric repulsion can counterbalance the gravitational collapse of a spherical body to achieve an equilibrium configuration. Stettner [39] showed that a spherical stellar structure in presence of net surface charge is found to be more stable than an uncharged one having a uniform density. Krasinski [40] demonstrated that preventing the gravitational collapse of spherically distributed matter into a point singularity necessitates the presence of a net charge. Through a generalisation of the Oppenheimer–Volkoff equation, Bekenstein [41] conducted a qualitative assessment of the stability of spheres composed of charged fluid. Over the years, several researchers have addressed the impact of charge on compact stellar structures [42, 43, 44, 45, 46, 47, 48, 49, 50].
@@
Recently, the $f(Q)$ theory has garnered a significant amount of attention from researchers worldwide. In the astrophysical scenario, it has provided new insights into compact stellar structures. In the context of charged configurations, Kaur et al. [51] presented a charged $f(Q)$ solution to study an anisotropic fluid distribution using the Vaidya-Tikekar metric ansatz. Maurya et al. [52] used an interesting approach of gravitational decoupling into the $f(Q)$ theory to study the constraining factors of mass and radii of the lighter component of GW190814 and other self-bound strange stars. Errehymy et al. [53] studied a singularity-free charged strange star model in the framework of $f(Q)$ gravity. Furthermore, Lohakare et al. [54] investigated the effects of gravitational decoupling on the maximum mass and stability of charged strange stars. In addition, several authors have included $f(Q)$ theory in their analysis of a stable stellar structure [55, 56, 57, 58].
Compact objects are extremely dense, meaning they have incredibly strong gravitational fields. The introduction of charge into the system introduces an additional repulsive force due to electrostatic Coulomb interaction. The specific form of the charge distribution helps ensure that the object remains stable under the interplay of these forces balancing gravitational attraction with the repulsion caused by the charge. A uniform or non-uniform charge distribution can influence the overall stability, so the form must be chosen carefully to ensure physical plausibility. The charge distribution, when introduced, must be compatible with the EoS to avoid violations of the energy conditions and to ensure that the object behaves consistently with relativistic models. If the charge is distributed too densely in certain regions, it can lead to unphysical behaviours, such as superluminal sound speeds or collapse under excess electrostatic pressure.The inclusion of charge modifies the spacetime geometry around the compact object. The Einstein-Maxwell field equations govern how the gravitational field interacts with the electromagnetic field generated by the charge. A specific form of charge ensures that the resulting geometry remains consistent with observed astrophysical phenomena. For example, a too concentrated charge could lead to the formation of horizons or singularities, which might be inconsistent with the object’s intended physical properties. Keeping all these points in mind and following the work of Gonçalves and Lazzari [59], we have chosen a specific form of charge distribution, $q=q_{0}r^{3}$ in this present manuscript which ensures a compatible relativistic solution. Additionally, it must be noted that the chosen ansatz for the electric charge distribution represents the simplest model commonly applied in charge distribution studies.
@@
In the present manuscript, the impact of charge has been investigated on the structure of a singularity-free compact star model within the framework of linear $f(Q)$ action. The advantage of $f(Q)$ gravity lies in its ability to derive the equations of GR without the necessity of employing the affine connection, thereby enhancing the inertial gravitational acceleration. This intriguing aspect coupled with the charged configuration reveals interesting stellar properties. Solving TOV equations, we have predicted the possible maximum mass and radius of compact stars in the framework of $f(Q)$ gravity in the presence of charge and we note some interesting results.
The paper is structured as follows: Section 2 demonstrates the fundamental ground work of $f(Q)$ theory in the presence of charge. Section 3 addresses the formulation and solution of the Einstein-Maxwell equations in the framework of $f(Q)$ gravity. The analytical expressions of the thermodynamic quantities like energy density $(\rho)$ and pressures (radial $(p_{r}$) and tangential $(p_{t})$) are obtained in this section. The effects of $f(Q)$ parameters on the boundary conditions are described in Section 4. Section 5 deals with the physical characteristic features of the model along with the determination of the maximum mass and radius, causality condition and energy conditions. The necessary stability criteria are addressed in Section 6. Finally, we conclude by summarising the major findings of the paper in Section 7.
@@ 2 Fundamentals of $f(Q)$ gravity in the presence of charge
where, $g$ defines the determinant of the fundamental metric tensor $g_{ij}$, i.e., $g=|g_{ij}|$, $f(Q)$ is a function of $Q$ known as non-metricity, $\lambda^{kij}_{l}$ and $\tau^{ij}_{k}$ are two Lagrangian multipliers, $R^{l}_{kij}$ and $T^{k}_{ij}$ are the Riemann tensor and torsion tensor, respectively, and $\mathfrak{L_{m}}$ and $\mathfrak{L_{q}}$ describe the Lagrangian density of matter and electromagnetic charge, respectively. Using affine connections $(\Gamma^{k}_{ij})$, the non-metricity is expressed as:
@@
    In Eq. (3), the term $L^{k}_{ij}$ represents the deformation and is expressed as:
@@
In absence of torsion as well as curvature, the affine connection can be effectively characterised by a tractable set of functions described as:
@@
Now, we obtain the tractable set of gravitational field equations by varying the Einstein-Hilbert action as written in Eq. (1) in terms of the metric tensor $g_{ij}$ of the form given below:
@@
Moreover, the variation of action given in Eq. (1) with respect to the affine connection yields,